# Generalized Linear Bandits with Limited Adaptivity

### v3

@@ B.3 Some Useful Lemmas
Let $\{x_{s}\}_{s=1}^{t}$ be a set of vectors. Define the sequence $\{\mathbf{V}_{s}\}_{s=1}^{t}$ as $\mathbf{V}_{1}=\lambda{\mathbf{I}}$, $\mathbf{V}_{s+1}=\mathbf{V}_{s}+x_{s}x_{s}^{\intercal}$ for $s\in[t-1]$. Further, let $\left\lVert x_{s}\right\rVert_{2}\leq L\ \forall\ s\in[t]$. Define the set $\{1=\tau_{1},\tau_{2}\ldots\tau_{m}=t\}$ such that: $\det(\mathbf{V}_{\tau_{i+1}})\geq 2\det(\mathbf{V}_{\tau_{i}})$ but $\det(\mathbf{V}_{\tau_{i+1}-1})<2\det(\mathbf{V}_{\tau_{i}})$ for $i\in\{2,\ldots m-1\}$. Then, the number of time doubling happens,i.e.$m$ is at most $O(d\log(t))$.
@@ Appendix C Useful Properties of GLMs
Putting $z=z_{2}$ establishes 24. To show 23, we further set $z=z_{1}+u(z_{2}-z_{1})$ for $u\in[0,1]$, (note that $z\geq z_{1}$) and integrate on $u$,
@@
For 25, we have, by application of 24, $\dot{\mu}(z_{1}+v(z_{2}-z_{1}))\geq\dot{\mu}(z_{1})\exp(R|v(z_{2}-z_{1})|)$. Therefore,
@@
$\geq\dot{\mu}(z_{1})\cdot\frac{1}{2+R|z_{1}-z_{2}|}$ (Lemma 10 of [Abeille et al., 2021])
@@
In [Faury, 2021] Section 1.4.1, the author claims that if the GLM parameter $z$ lies in a bounded set, then the GLM is self-concordant, i.e., $|\ddot{\mu}(z)|\leq a\dot{\mu}(z)$, for some appropriate constant $a$ over this bounded set. Thereafter the author notes that the techniques developed in [Faury, 2021] guarantees $\kappa$-free regret rates (in $\sqrt{T}$ term) for such GLMs (i.e., all GLMs with bounded parameter). However, the claim regarding self-concordance of GLMs is not true in general. There are classes of GLMs whose parameters may be restricted in a bounded set, but for them no constant $a$ exists. One such example is the exponential distribution. The link function $\mu$ for exponential distribution is given as $\mu(z)=-\frac{1}{z}$. If we allow $z$ to lie in the set $(-c,0)$ for some positive $c$, then we have $\mu(z)$ strictly increasing (satisfying our assumption on monotonicity of $\mu$, thus a valid example). However, for this GLM,
@@ Appendix D Computational Cost
Consider a log-loss minimization oracle that returns the unconstrained MLE for a given GLM class with a computational complexity of $C_{opt}\cdot n$, when the log-loss is computed over $n$ data points. Let the maximum number of arms available every round be $K$. Furher, let the computational cost of an oracle that solves the non-convex optimization 6 be $NC_{opt}$.
@@
Computational Cost of RS-GLinCB: In the RS-GLinCB algorithm, the estimator $\widehat{\theta}_{w}$ is computed during each warmup round. Additionally, during non-warmup rounds, the estimator $\widehat{\theta}$ is computed a maximum of $O(\log(T))$ times. These computations involve utilizing both the log-loss oracle and the non-convex projection oracle. Furthermore, in each non-warmup round, the algorithm executes an elimination step. This yields an amortized time complexity of $O(C_{opt}\log{T}+NC_{opt}\log^{2}(T)+K)$ per round.
@@
Comparison with ECOLog [Faury et al., 2022] shows that execution time for RS-GLinCB is significantly smaller. We posit that this is because RS-GLinCB solves a large convex optimization problem but less frequently, resulting into smaller overhead at the implementation level, while ECOLog solves a smaller convex optimization problem, but does so every round. On an implementation level, this translates into more function calls and computation. Further, we observe that with increasing $\kappa$, the execution time of RS-GLinCB increases, which is in accordance with Lemma 4.1 that quantifies the number of policy switches as an increasing function of $\kappa$.
While comparing with GLOC [Jun et al., 2017], we observe that RS-GLinCB performs better than GLOC in both high and low $\kappa$ regimes. Since GLOC runs an online convex optimization (online Newton step) algorithm to generate its confidence sets, the time taken by GLOC is nearly constant with changing $\kappa$. On the other hand, in accordance with Lemma 4.1, the computational cost of RS-GLinCB increases with $\kappa$. However, after a few initial rounds, when neither of the switching criteria are triggered, RS-GLinCB does not need to solve any computationally intensive optimization problem, hence these rounds execute very fast. In practice, with typical data distribution, RS-GLinCB reaches this stage much before what the worst-case guarantees show, hence we see it perform better than GLOC.
@@ E.1 Convex Relaxation
We now follow closely the proof outlined in Appendix B.3 of [Abeille et al., 2021] with minor changes. By second-order Taylor’s expansion, for any $\theta\in\mathbb{R}^{d}$, we can write
@@
Recall that $\widehat{\theta}$ is the unconstrained MLE, therefore $\nabla{\mathcal{L}}_{t}(\widehat{\theta})=\mathbf{0}$. By a similar Taylor expansion as above and some algebraic manipulations (see Appendix B.3 of [Abeille et al., 2021]), we have, for $\theta^{*}$.

### v4

@@ B.6 Some Useful Lemmas
Let $\{x_{s}\}_{s=1}^{t}$ be a set of vectors. Define the sequence $\{\mathbf{V}_{s}\}_{s=1}^{t}$ as $\mathbf{V}_{1}=\lambda{\mathbf{I}}$, $\mathbf{V}_{s+1}=\mathbf{V}_{s}+x_{s}x_{s}^{\intercal}$ for $s\in[t-1]$. Further, let $\left\lVert x_{s}\right\rVert_{2}\leq L\ \forall\ s\in[t]$. Define the set $\{1=\tau_{1},\tau_{2}\ldots\tau_{m}=t\}$ such that: $\det(\mathbf{V}_{\tau_{i+1}})\geq 2\det(\mathbf{V}_{\tau_{i}})$ but $\det(\mathbf{V}_{\tau_{i+1}-1})<2\det(\mathbf{V}_{\tau_{i}})$ for $i\in\{2,\ldots m-1\}$. Then, the number of time doubling happens,i.e.$m$, is at most $O(d\log(t))$.
@@ Appendix C Useful Properties of GLMs
Putting $z=z_{2}$ establishes 26. To show 25, we further set $z=z_{1}+u(z_{2}-z_{1})$ for $u\in[0,1]$, (note that $z\geq z_{1}$) and integrate on $u$,
@@
For 27, we have, by application of 26, $\dot{\mu}(z_{1}+v(z_{2}-z_{1}))\geq\dot{\mu}(z_{1})\exp(R|v(z_{2}-z_{1})|)$. Therefore,
@@
$\geq\dot{\mu}(z_{1})\cdot\frac{1}{2+R|z_{1}-z_{2}|}$ (Lemma 10 of Abeille et al. [2021])
@@
In Faury [2021] Section 1.4.1, the author claims that if the GLM parameter $z$ lies in a bounded set, then the GLM is self-concordant, i.e., $|\ddot{\mu}(z)|\leq a\dot{\mu}(z)$, for some appropriate constant $a$ over this bounded set. Thereafter the author notes that the techniques developed in Faury [2021] guarantees $\kappa$-free regret rates (in $\sqrt{T}$ term) for such GLMs (i.e., all GLMs with bounded parameter). However, the claim regarding self-concordance of GLMs is not true in general. There are classes of GLMs whose parameters may be restricted in a bounded set, but for them no constant $a$ exists. One such example is the exponential distribution. The link function $\mu$ for exponential distribution is given as $\mu(z)=-\frac{1}{z}$. If we allow $z$ to lie in the set $(-c,0)$ for some positive $c$, then we have $\mu(z)$ strictly increasing (satisfying our assumption on monotonicity of $\mu$, thus a valid example). However, for this GLM,
@@ Appendix D Computational Cost
Consider a log-loss minimization oracle that returns the unconstrained MLE for a given GLM class with a computational complexity of $C_{opt}\cdot n$, when the log-loss is computed over $n$ data points. Let the maximum number of arms available every round be $K$. Furher, let the computational cost of an oracle that solves the non-convex optimization 8 be $NC_{opt}$.
@@
Computational Cost of RS-GLinCB: In the RS-GLinCB algorithm, the estimator $\widehat{\theta}_{o}$ is computed each time Switching Criterion I is triggered. Additionally, during rounds when Switching Criterion I is not triggered, the estimator $\widehat{\theta}_{\tau}$ is computed a maximum of $O(\log(T))$ times. These computations involve utilizing both the log-loss oracle and the non-convex projection oracle. Furthermore, in each non-Switching Criterion I round, the algorithm executes an elimination step. This yields an amortized time complexity of $O(C_{opt}\log{(T)}+NC_{opt}\log^{2}(T)+K)$ per round.
@@
Comparison with ECOLog Faury et al. [2022] shows that execution time for RS-GLinCB is significantly smaller. We posit that this is because RS-GLinCB solves a large convex optimization problem but less frequently, resulting into smaller overhead at the implementation level, while ECOLog solves a smaller convex optimization problem, but does so every round. On an implementation level, this translates into more function calls and computation. Further, we observe that with increasing $\kappa$, the execution time of RS-GLinCB increases, which is in accordance with Lemma 4.1 that quantifies the number of policy switches as an increasing function of $\kappa$.
While comparing with GLOC Jun et al. [2017], we observe that RS-GLinCB performs better than GLOC in both high and low $\kappa$ regimes. Since GLOC runs an online convex optimization (online Newton step) algorithm to generate its confidence sets, the time taken by GLOC is nearly constant with changing $\kappa$. On the other hand, in accordance with Lemma 4.1, the computational cost of RS-GLinCB increases with $\kappa$. However, after a few initial rounds, when neither of the switching criteria are triggered, RS-GLinCB does not need to solve any computationally intensive optimization problem, hence these rounds execute very fast. In practice, with typical data distribution, RS-GLinCB reaches this stage much before what the worst-case guarantees show, hence we see it perform better than GLOC.
@@ E.1 Convex Relaxation
We now follow closely the proof outlined in Appendix B.3 of Abeille et al. [2021] with minor changes. By second-order Taylor’s expansion, for any $\theta\in\mathbb{R}^{d}$, we can write
@@
Recall that $\widehat{\theta}$ is the unconstrained MLE, therefore $\nabla{\mathcal{L}}_{t}(\widehat{\theta})=\mathbf{0}$. By a similar Taylor expansion as above and some algebraic manipulations (see Appendix B.3 of Abeille et al. [2021]), we have, for $\theta^{*}$.